Green Thumb Gardening is a small gardening service that uses activity-based costing to estimate costs for pricing and other purposes. The proprietor of the company believes that costs are driven primaily by the size of customer lawns, the size of customer garden beds, the distance to travel to customers, and the number of customers. In addition, the costs of maintaining garden beds depends on whether the beds are low maintenance beds (mainly ordinary trees and shrubs) or high maintenance beds (mainly flowers and exotic plants). Accordingly, the company uses the five activity cost pcols listed below Activity Cost Pool Caring for lawn Caring for garden beds-low maintenance Caring for garden beds-high maintenance Travel to jobs Customer billing and service Activity Measure Square feet of lawn Square feet of low maintenance beds Square feet of high maintenance beds Miles Number of customers The company has already completed its first stage allocations of costs and has summarized its annual costs and activity as follows: Estimated Activity Cost Pool Caring for lawn Caring for garden beds-low maintenance Caring for garden beds-high maintenance Travel to jobs Customer billing and service Cost S 83,400 S 36,000 S 54,200 $ 5,200 S 10,700 Expected Activity 190,000 square feet of lawn 26,000 square feet of low maintenance beds 20,000 square feet of high maintenance beds 17,000 miles 28 custorners Required Compute the activity rate for each of the activity cost pools. (Round your answers to 2 decimal places.) Cost Pool Activity Rate Caring for lawn Caring for garden beds low maintenance Caring for garden beds high per square foot of lawn per square foot of low maintenance beds per square foot of high maintenance maintenance Travel to jobs Customer billing and service beds per mile per customer
Answer #1

Caring for lawn = $83,400 / 190,000 = $0.44 per square foot of lawn

Caring for garden beds-low maintenance = $36,000 / 26,000 = $1.38 per square foot of low maintenance beds

Caring for garden beds-high maintenance = $54,200 / 20,000 = $2.71 per square foot of high maintenance beds

Travel to jobs = $5,200 / 17,000 = $0.31 per mile

Customer billing and service = $10,700 / 28 = $382.14 per customer